CAT-6 twisted pair cable is a commonly used computer networks cable that provides a frequency bandwidth of up to 250 MHz. It is suitable for most varieties of Ethernet over twisted pair up to 1000 BASE-T (Gigabit Ethernet). Determine whether a signal-to-noise ratio (SNR) of 20 dB is adequate to transmit 0.9 Gbps on this cable?

1
Expert's answer
2022-02-03T08:39:14-0500

CAT-6 twisted pair cable is a commonly used computer networks cable




SNR(dB) =10log10(SNR)SNR(dB)=10log10(SNR)


Therefore:


20=10log10(SNR)


SNR=102 =>100



To get Capacity:


C=Blog2(1+SNR)C=Blog2(1+SNR)


C=0.9*109log2(101)


C=5 992.39MHz


Hence the signal to noise ratio of 20dB is inadequate